Built-Rite Corp is evaluating an extra dividend versus a share repurchase. In either case $20,000 would be spent. Current earnings are $6.0 per share, and the stock currently sells for $40 per share. There are 4,000 shares outstanding. Ignore taxes and other imperfections in answering parts (a) and (b).

a.  Evaluate the two alternatives in terms of the effect on the price per share of the stock and shareholder wealth. 

b.  What will be the effect on Built-Rite Corp’s EPS and PE ratio under the two different scenarios? 

c.  In the real world, which of these actions would you recommend? Why?  



1
Expert's answer
2022-02-09T10:43:46-0500

A)

Alternative 1, extra Dividends:

Dividends per share =20000 ÷ 4000

= 5

Price of stock after dividends = 40 - 5

= 35

Share holders wealth per share = 35+5

= 40

Alternative 2 Repurchases:

Prices per share = 40

Share holders wealth per share at market price or cash (if he agrees to repurchase)

= 40

B)

Alternative 1, extra dividends:

EPS = 6

PE ratio = Market price ÷ Earning per share

= 35÷6

= 5.8333

Alternative 2, repurchases:

EPS = Total earnings ÷ current number of shares

= (6 × 4000) ÷ (4000 - (20000÷40))

= 24000 ÷ 3500

= 6.8571

PE ratio = Market price ÷ earning share

= 40÷ 6.8571

= 5.8334

c)

Recommendation:

The sales of shares have the best chance of success. However, Built-Rite Corp can boost their EPS figures through stock buybacks that reduce the number of outstanding shares. So, strong profit growth also demands strong sales growth.